Minden Company introduced a new product last year for which it is trying to find an optimal selling price. Marketing studies suggest that the company can increase sales by 5,000 units for each $2 reduction in the selling price. The company’s present selling price is $92 per unit, and variable expenses are $62 per unit. Fixed expenses are $839,400 per year. The present annual sales volume (at the $92 selling price) is 25,500 units.

Required:
1.

What is the present yearly net operating income or loss?

  

2.

What is the present break-even point in unit sales and in dollar sales?

  

3.

Assuming that the marketing studies are correct, what is the maximum annual profit that the company can earn? At how many units and at what selling price per unit would the company generate this profit?

  

4.

What would be the break-even point in unit sales and in dollar sales using the selling price you determined in (3) above (e.g., the selling price at the level of maximum profits)?
Answer #1

1) Calculate present net income (loss)

Sales (25500*92) 2346000
Variable expense 1581000
Contribution margin 765000
Fixed cost 839400
Net income (loss) -74400

2) Break even point = Fixed cost/Contribution margin per unit = 839400/30 = 27980 Units

Break even sales = 27980*92 = 2574160

3) Calculate following

Selling price Unit Profit
92 25500 -74400
90 30500 14600
88 35500 83600
86 40500 132600
84 45500 161600
82 50500 170600
80 55500 -159600

So maximum selling price per unit = $82 and unit sold = 50500 Units and general profit = 170600

4) Break even unit = 839400/20 = 41970 Units

Break even sales = 41970*82 = $3441540